I don't know much about how the cron job runs, I'll forward this to Max.
I started to receive emails from cron every 15min. Is the email related to this? And is it normal? I never received these cron emails before when the sum-page was running.
Max says it should be fixed now. Have the emails stopped?
Circuit1: It is nice to receive the voltage across the transimpedance resistor with a high impedance buffer (or amplifier), as close to the resister as possible. This amplifier needs to have low numbers for input bias current, input offset current, and input current noise. These current noise becomes the noise of the temperature reading. On the top of that, the input voltage noise of the buffer will be added to the output. The typical noise model can be found in http://www.analog.com/media/en/technical-documentation/application-notes/AN-940.pdf
The good candidates for the buffer is LT1128, ADA4004, OPA140, and LT1012. If the application is not too sensitive to the total noise, OPA604 is a good choise with easier handling.
Circuit2: With the same reason, AD741 is an old generic amp that is not a great choise for this purpose. The current noise is more significant because of the higher transimpedance here. The same noise model as above can be used to analyze the performance.
I wanted to know what this Vmon exactly is. D010001 is telling us that the Vmon channels are HPFed with fc=30Hz (Attachment 1). Is this true?
I checked the quiscent noise spectrum of the ITMX UL coil output (C1:SUS-ITMX_ULCOIL_OUT) and the corresponding VMON (C1:SUS-ITMX_ULVmon). (Attachment 2 Ref curves). I did not find any good coherence. So the nominal quiscent Vmon output is carrying no useful information.
Question: How much do we need to excite the coil output in order to see any meaningful signal?
As I excite the ITMX UL coil (C1:SUS-ITMX_ULCOIL_EXC) with uniform noise of 100-300 counts below 0.3Hz, I eventually could see the increase of the power spectrum and the coherence (Attachment 2). Below 0.1 Hz the coherence was ~1 and the transfer function was measured to be -75dB and flat. But wait, why is the transfer function flat?
In fact, if I inject broadband noise to the coil, I could increase the coil output and Vmon at the same time without gaining the coherence. (Attachment 3). After some more investigation, I suspect that this HPF is diabled (= bypassed) and aliasing of the high freq signal is causing the noise in Vmon.
In order to check this hypothesis, we need to visit the board.
The main C1 summary pages are back online now thanks to Max and Duncan, with a gap in pages from June 8th to July 4th. Also, I've added my new VMon and Sensors tabs to the SUS parent tab on the main pages. These new tabs are now up and running on the July 7th summary page.
Here's a link to the main nodus pages with the new tabs: https://nodus.ligo.caltech.edu:30889/detcharsummary/day/20160707/sus/vmon/
And another to my ldas page with the tabs implemented: https://ldas-jobs.ligo.caltech.edu/~praful.vasireddy/1150848017-1150848317/sus/vmon/
Let me know if you have any suggestions or see anything wrong with these additions, I'm still working on getting the scales to be right for all graphs.
I am using AD592CNZ temperature transducer ICs for measuring temperature inside as well as outside the enclosure. It is a current output IC which outputs current proportional to temperature. As mentioned in the data sheet of AD592, I am using the following two schematics:
Though I still need to calibrate these temperature transducers, I did some measurements. I have temperature readings, and now my goal in few days is to find a transfer function of temperature fluctuations inside the enclosure to outside the enclosure.
About data acquisition:
We have re-configured the raspberry pi(B8:27:EB:70:D0:D8) on martian network. It's new ip address is 192.168.113.107(domenica.martian). Also, we have added the Acromag Busworks card(00:01:C3:00:9F:C8) on the martian network and its ip address is 192.168.113.237(acroey.martian).
I've added a new tab for VMon under the SUS parent tab. I'm still working out the scale and units, but let me know if you think this is a useful addition. Here's a link to my summary page that has this tab: https://ldas-jobs.ligo.caltech.edu/~praful.vasireddy/1151193617-1151193917/sus/vmon/
I'll have another tab with VMon BLRMS up soon.
Also, the main summary pages should be back online soon after Max fixed a bug. I'll try to add the SUS/VMon tab to the main pages as well.
The DAFI block was reviewed by Rana yesterday. The following changes/improvements were suggested: (Updated on 20th July 2016 with tasks taat remain in red)
1) include all the various channels like PEM, LSC, ASC, SUS, SEI, etc. as the inputs. Currently the inputs are only the LSC.
2) include all the control signals.
3) create a very detailed diagram of the entire signal flow and plan tasks accordingly.
4) Enable cascading of various DSP processes.
5) Adjusting the gain of the AGC such that the amplitude of the output signal comes to about half the peak amplitude offered by the ADC. This will help taking advantage of the entire dynamic range of the ADC.
6) change the enable button styles from a text input based controller to a button controller.
7) Currently, disabling a particular signal terminates the signal. Instead, it should turn into a unity gain block on disabling.
8) Check if the Fibox does AC coupling or not. If not, add an AC coupling arrangement in the DAFI.
9) Check the nature of the ADC1 and ADC2 inputs to the DAFI. I checked them yesterday, and they are channels 25 and 26 of ADC0, which are empty.
Vacuum Status: Chamber Open
All chamber annuloses are vented. Vac Monitor screen is not communicating with gauges. The valve position indicator are working.
RGA is pumped by Maglev through VM2
Yesterday Q noticed that PRM_sensor_LR was 0.098V This actually went to ~ zero on 7-3
The last RGA scan of this pumpdown 78
Pressure plot of 640 days long pd 78
CC1 cold cathode gauge was jump started with an accidental pressure glitch, that you can see on P1 plot
With Koji's help, I've hacked together an arrangement that will allow us to monitor the output of the coil driver to the UL coil.
The arrangement consists of a short custom ribbon cable with female DB25 connectors on both ends - the particular wire sending the signal to the UL coil has a 100 ohm resistor wired in series, because the coil has resistance ~20ohm, and the output of the coil driver board has a series 200(?) ohm resistor, so by directly monitoring the voltage at this point, we may not see a glitch as it may register too small. Tangentially related: the schematic of the coil driver board suggests that the buffered output monitor has a gain of 0.5.
To monitor the voltage, I use the board to which the 4 Oplev signals are currently hooked up. Channel 7 on this particular board (corresponding to ADC channel 30 on c1scx) was conveniently wired up for some prior test, so I used this channel. Then, I modified the C1SCX model to add a testpoint to monitor the output of this ADC. Then, I turned OFF the input on the coil output filter for the UL Coil (i.e. C1:SUS-ETMX_ULCOIL_SW1) so that we can send a known, controlled signal to the UL Coil by means of awggui. Next, I added an excitation at 5 Hz, amplitude 20 counts (as the signal to the coil under normal conditions was approximately of this amplitude) to the excitation channel of the same filter module, which is the state I am leaving the setup in for the night. I have confirmed that I see this 5Hz oscillation on the monitor channel I set up. Oddly, the 0 crossings of the oscillations happen at approximately -1000 counts and not at 0 counts. I wonder where this offset is coming from? The two points I am monitoring the voltage across is shown in the attached photograph - the black clip is connected to the lead carrying the return signal from the coil.
I also wanted to set up a math block in the model itself that monitors, in addition to the raw ADC channel, a copy from which the known applied signal has been cancelled, as presumably a glitch would be more obvious in such a record. However, I was unable to access the excitation channel to the ULCOIL filter from within the SCX model. So I am just recording the raw output for tonight...
One glitch was seen to occur without a change in the output voltage monitors in ELOG 11744
I'd suggest clamping and moving it to the flow bench so you can inspect with a bright light. Then remove the wire and inspect the standoff, but hurry up with getting it in the soak bath so you can start on the cleaning of the other ones.
I wonder if we're really sure that its a mechanical problem with ETMX.
Gautam tells me that the local damping was always ON when looking for the jumps. This means that the coil driver was still hooked up and we can't rule out glitches in the DAC or the coil driver.
The UL OSEM shows the biggest movement (10 microns). The LR shows the second most (6-7 microns). The others are 2x less. So its consistent with a voltage change on UL,
Is this consistent with a slip in one of the wire standoffs? I think no.
After hardware errors prevented me from using optimus, I switched my generation of summary pages back to the clusters. A day's worth of data is still too much to process using one computer, but I have successfully made summary pages for a timescales of a couple of hours on this site: https://ldas-jobs.ligo.caltech.edu/~praful.vasireddy/
Currently, I'm working on learning the current plot-generation code so that it can eventually be modified to include an interactive component (e.g., hovering over a point on a timeseries would display the GPS time). Also, the 40m summary pages have been down for the past 3 weeks but should be up and working soon as the clusters are now alive.
Rough summary of today's progress:
I didn't really see anything out of the ordinary on the ETMX suspension. Earthquake stops had clearance, OSEMS were secure, no visible glue degredation on face magnets. Inspection with green LED flashlight didn't reveal any obscene dirtieness on either face, just a few particles here and there. The top of the opic barrel unsurprisingly has a good amount of particulate. The wire grooves are way too small to resolve anything at this point, other than that they exist.
The suspension footprint is already marked, tomorrow we can move the suspension closer to the door to get an even closer look at it, before removing it from the chamber.
I am trying to design an antialiasing filter, which also has two switchable whitening stages. I have designed a first version of a PCB for this.
The board takes differential input through PCB mountable BNCs. It consists of an instrumentaiton amplifier made using quad opamp ADA4004, followed by two whitening blocks, also made using ADA4004, which can be bypassed if needed, depending upon a control input. The mux used for this purpose is Maxim MAX4158EUA. These two whitening blocks are followed by 2 the LPF stages. A third LPF stage could be added if needed. These use AD829 opamps. After the LPFs are two amplifiers for giving a differential output through two output BNCs. The schematic is shown in attachment 1: "AA.pdf". The top layers of the layout are shown in attachment 2 (AAtop.pdf), the bottom layers in attachment 3 (AAbottom.pdf), and the entire layout in attachment 4 (AAbrd.pdf).
The board has 6 layers (in the order from top to bottom):
1) Top signal layer;
2) Internal plane 1 (GND),
3) Internal plane 2 (+15V),
4) Internal plane 3 (-15V),
5) Internal plane 4 (GND),
6) Bottom signal layer.
Power: +15, -15 and GND is given through a 4 pin header connector.
The dimensions of the board are 1550 mil 6115 mil (38.1mm155.3mm) and the overall dimensions including the protruding BNC edges are 1550 mil 7675 mil (38.1mm194.9mm)
I would like to have inputs on the layout telling me if any component/trace needs to be changed/better placed, any other things about the board need to be changed, etc.
P.S.: I have also added a zipped folder "AA.zip" containing the schematic and board files, as well as the above pdfs.
Glass soaking dish with teflon guides.
Proposed Acetone soak dish for SOS epoxy softening.
It has good acces through 5" top ID. The set up is stable and teflon lined.
Materials: glass jar with SS cover, teflon bricks, 0.008" teflon wrapped "high density Drever bricks" and aluminum
Drever brick: I beleive it is a Tungsten alloy. We used it as vac-bat savor at the coffe can. It has high density, heavy and hard, it was never identified.
I will soak one brick to see if it has any reaction ability with acetone.
NO means that only Glass and Teflon can be used for this fixture in Acetone. We can not take a chance on the coating!
I guess the small surface area Aluminum dumbbell, guide rod and-or wire standoff, magnet and epoxy does not degrade the acetone such way that it effects our coating.
Not ot mention, that only the very edge of the coating would in this solution.
Here are some plans / rough procedures for this week's vent. It is unlikely that I have though of everything, but this should be a reasonable starting point.
The mode cleaner still hasn't been locked in air, we may not want to touch the Y arm optics until we are able to lock to the Y arm and dither align, so we are sure to keep the input pointing from drifting away too much.
For $optic in [ITMX, ITMY, ETMY]:
There has been an ongoing memory error in optimus with the following messages:
Message from syslogd@optimus at Jun 30 14:57:48 ...
kernel:[1292439.705127] [Hardware Error]: Corrected error, no action required.
Message from syslogd@optimus at Jun 30 14:57:48 ...
kernel:[1292439.705174] [Hardware Error]: CPU:24 (10:4:2) MC4_STATUS[Over|CE|MiscV|-|AddrV|CECC]: 0xdc04410032080a13
Message from syslogd@optimus at Jun 30 14:57:48 ...
kernel:[1292439.705237] [Hardware Error]: MC4_ADDR: 0x0000001ad2bd06d0
Message from syslogd@optimus at Jun 30 14:57:48 ...
kernel:[1292439.705264] [Hardware Error]: MC4 Error (node 6): DRAM ECC error detected on the NB.
Message from syslogd@optimus at Jun 30 14:57:48 ...
kernel:[1292439.705323] [Hardware Error]: cache level: L3/GEN, mem/io: MEM, mem-tx: RD, part-proc: RES (no timeout)
Optimus is a Sun Fire X4600 M2 Split-Plane server. Based on this message, the issue seems to be in memory controller (MC) 6, chip set row (csrow) 7, channel 0. I got this same result again after installing edac-utils and running edac-util -v, which gave me:
mc6: csrow7: mc#6csrow#7channel#0: 287 Corrected Errors
and said that all other DIMMs were working fine with 0 errors. Each MC has 4 csrows numbered 4-7. I shut off optimus and checked inside and found that it consists of 8 CPU slots lined up horizontally, each with 4 DIMMs stacked vertically and 4 empty DIMM slots beneath. I'm thinking that each of the 8 CPU slots has its own memory controller (0-7) and that the csrow corresponds to the position in the vertical stack, with csrow 7 being the topmost DIMM in the stack. This would mean that MC 6, csrow 7 would be the 7th memory controller, topmost DIMM. The channel would then correspond to which one of the DIMMs in the pair is faulty although if the DIMM was replaced, both channels 0 and 1 would be switched out. Here are some sources that I used:
I'll find the exact part needed to replace soon.
I just disconnected the 6th instrument grade air cylinder from the vacuum envelope at 720 Torr. Now it will reach equilibrium through a filter as it sucks in lab air.
This is the sure way not to over pressurize the chamber.
1, Fix ETMX sus "jump issue"
2, First Contact clean the arms
3, Install new spare cold cathode and convectron gauges: InstruTech-Hornet
4, Install 50 mm apeture beam baffles
5, Check and clean optical quality viewport from inside
The following bullets were executed:
Check crane functionality & cleanliness last week
We are venting the 40m IFO
Steve has ordered some teflon parts to take the place of the metal parts in his acetone-soaking jig. They should arrive tomorrow.
So, we will be begin the venting process tomorrow. Doors to come off on Tuesday.
I have transferred most of the temperature measurement stuff from the front area to seismometer at the end of Y-arm. While arranging the components I have taken all care that they will not interfere with existing system. Also, I have temporarily taken a monitor from the front area to the area near same seismometer as I couldn't talk to Rpi via ssh. For next twelve hours, I am now recording temperature inside as well as outside the seismometer enclosure. Some temperature sensors are inside the enclosure while some are outside the seismometer enclosure.
Last night, we set about trying to see if we could measure and verify the predictions of the simulations, and if there are indeed HOM sidebands co-resonating with the carrier. Koji pointed out that if we clip the transmitted beam from the arm incident on a PD, then the power of the higher order HG modes no longer integrate to 0 (i.e. the orthogonality is broken), and so if there are indeed some co-resonating modes, we should be able to see the beat between them on a spectrum analyzer. The procedure we followed was:
We then repeated the above steps at the X-end (but here, an additional lens had to be installed to focus the IR beam onto the PDA10CF - there was, however, sufficient space on the table so we didn't need to remove the PDA520 for this measurement).
Y-end: DC power on the photodiode at optimal alignment ~ 200mV => spectra taken by deliberately misaligning the beam incident on the PD till the DC power was ~120mV (see remarks about these values).
I converted the peak heights seen on the spectrum analyzer in volts to power by dividing by transimpedance (=5*10^3 V/A into a 50ohm load) * responsivity at 1064nm (~0.6A/W for PDA10CF).
I have updated the vent prep checklist on the wiki. Gautam and I did the following things from it:
Reduce input power to no more than 100mW by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.)
The following bullets have not yet been executed:
Check crane functionality & cleanliness
I've gone through the SOS suspension document (E970037) and some old elogs to get an idea of all the accesories we need for the process of suspending, aside from the tower itself, which Steve has already put together. Gautam and I have laid our eyes upon most of the critical pieces. Some other objects are unknown, and perhaps not strictly neccesary.
Confirmed to exist:
In addition, I am told that we have a long ribbon cable that can run from the X end to the clean room to enable OSEM damping control while we do the pitch alignment.
Things mentioned in the procedure I have not found:
Some other tasks and their status:
In the attached photo from 2012, one can see the installed black glass baffle. According to the drawings (LIGO-DNNNXXX) this one has a clear aperture of 40 mm.
In (someplace ?) we have clean baffles with a 50 mm aperture which can be installed during this vent. In order to be more conservative, let us choose to swap these out for all 4 test masses during the upcoming vent using the green laser as an alignment guide, as Koji described at today's lunch meeting.
They are located at the top of E1 drawer cabinet
Praful Vasiceddy received 40m specific basic safety training.
Aakash Patil received 40m specific basic safety training.
Hello, I am Varun Kelkar. I will be working at the 40m lab as a SURF student this summer with Eric Quintero on Audio processing for real time control system signals. This week I will mostly be working on implementing basic DSP C-code offline. Currently I am trying to write a code for noise whitening.
Varun has received 40m specific basic safety training today.
Hi Steve -
I found the doc I was looking for:
Specifically, you might find guidance in Section 5 and the pictures at the end of the doc. This should work for Vacseal as well.
Good luck - it will take some time (hours to day or 2)...
I'll be interested to know how it goes.
GariLynn helped us develop this procedure so you could also ask her to cast an eye over the setup if you are worried.
ps: there is no existing fixture to hold SOS optic while soaking it
We'll follow LIGO policy:
Our policy is to use first contact within 1 year of purchase for use in the interferometers. For inspection use I am comfortable with out-of-date use.
GaryLinn offered their indate First Contact for use.
I have taken out the heaters and temperature sensors from the enclosure which was made by Megan last summer. Soon I will test and configure those heaters.
The existing enclosure for seismometer at LIGO 40m lab is a cylindrical stainless steel can placed upside down over the seismometer. It has more empty space between the seismometer and the internal surface of enclosure which is not desirable(I'll quantitatively elaborate this statement once my temperature measuring setup is ready).
Stainless steel has a thermal conductivity in the range of 16.3 to 16.7 W/m/K and magnetic permeability 1.260e-6 H/m.Assuming an ambient temperature 298K, and the temperature inside the enclosure as 295K, as well as substituting all the values for dimesions and material properties of existing enclosure,
k=16.4 W/mK, μ=1.260e-6 H/m, L=2ft=0.6096m, b=r2 =0.5ft=0.1524m, thickness=5mm, a=r1 =0.1474m.
So by using the textbook relations(I have mentioned them in my report), the value of attenuation coefficient is 5.953584e-05 and the value of rate of heat transfer= 5.64913 kW. The attenuation coefficient value is quite better for steel but proper care needs to be taken to avoid heat transfer. For studying the variation of rate of heat transfer and attenuation with the thickness of enclosure material, I have plotted the following attached graphs for different materials which include hardened stainless steel, aluminium, pure iron and nanoperm-muMetal.
About Data Acquisation
I have already invested a lot of time to configure and use acromag busworks card over ethernet. So now I have made an arrangement to measure temperature by AD592CNZ temperature transducer IC. I would be using raspberry pi for acquiring data untill I figure out a way to use acromag busworks card for the same. This setup of acquiring logging temperature using raspberry pi is mostly ready except the calibration part.
Made a dry run of the in-situ cleaning for a 3inch optic.
Attachment 1: The Al dummy mass is clamped in the suspension cage.
Attachment 2: The front surface was painted. The nominal brush with the FC bottle was used.
Attachment 3: Zoom in of the front surface.
Attachment 4: The back surface was painted.
Attachment 5: The back surface was peeled.
Attachment 6: The front surface was peeled too.
Attachment 7: The peeled layers.
1. To paint a thick layer (particlarly on the rim) is the key to peel it nicely.
2. It was helpful for easier peeling to have mutiple peek tabs. Two tabs were sufficient for ~1" circle.
3. The nominal brush with the bottle was OK although one has to apply the liquid many times to cover such a large area. A larger brush may cause dripping.
4. The nominal brush was sufficiently long once the OSEMs are removed. In any case it is better to remove the OSEMs.
Found 60 EM172 microphones. Previous elog with details: 7777.
Found 1 out of 2 bluebird microphones in the 40m.
Set up gwsumm on optimus and generated summary pages from both L1 and C1 data. Still a few manual steps need to be taken during generation, not fully automated due to some network/username issues. nds2 now working from optimus after restarting nds2 server.
Having investigated the mode-overlap as a function of RoC of the PRC and SRC folding mirrors, I've now been looking into possible stability issues, with the help of some code that EricQ wrote some time back for a similar investigation, but using Finesse to calculate the round trip Gouy phase and other relevant parameters for our current IFO configuration.
To do so, I've been using:
As a first check, I used flat folding mirrors to see what the HOM coupling structure into the IFO is like (the idea being then to track the positions of HOM resonances in terms of CARM offset as I sweep the RoC of the folding mirror).
However, just working with the flat folding mirror configuration suggests that there are order 2 22MHz and order 4 44MHz HOM resonances that are really close to the carrier resonance (see attached plots). This seems to be originating from the fact that the Y-arm length is 37.81m (while the "ideal" length is 37.795m), and also the fact that the ETM RoCs are ~3m larger than the design specification of 57m. Interestingly, this problem isn't completely mitigated if we use the ideal arm lengths, although the order 2 resonances do move further away from the carrier resonance, but are still around a CARM offset of +/- 2nm. If we use the design RoC for the ETMs of 57m, then the HOM resonances move completely off the scale of these plots...
Given the effect on the contrast defect, the consensus at the meeting Wednesday 6-22 was that we should continue to use the existing ETMX optic.
(on tower )
NOT finished, last edited 6-28
I got some QPR Nitrile gloves. They are LIGO approved.White nitrile gloves are naturally anti-static- 109 ohms
Their touch not as good as laytex gloves but try to use them.
The fire alarm came on around 15:05 for about 2-3 minutes. We all left the lab and counted heads. I called Paul Mackel x2646 (cell 626/ 890- 3259) at Fire Protection Services. He said that this alarm test was planned and we should of got an email notice. Perhaps I missed that notes.
Fire alarm went off several minutes ago. Talked to security and they said there was no fire. It beeped twice again just now. No one has been working on the IFO today.
Using an RC to BNC connector from the inner drawer, I have added a second output cable going from the output Fibox in the control room to the audio mixer.
I have updated the DAFI with the following changes:
1) Separated both the channels of stereo output completely, as well as in the GUI.
2) Added text monitors for the inputs and outputs.
The stereo output is now ready except for a cable going from the second channel of the output fibox to the audio mixer.
Attached is the main DAF_OVERVIEW screen and its link button from the LSC screen labelled "DAFI"
I wish to have stereo audio output for the DAF module. Hence, there needs to be a second output from the DAF. I added this second output to the model. Following are the details:
FiBox: It consists of two analog inputs which are digitized and multiplexed and transmitted optically. (only 1 fiber is needed due to multiplexing). Attachment 1 shows the fibox with its 2 analog inputs (one of which, is connected), and 1 fiber output. The output of the DAF goes to the FiBox. Until today, the Fibox recieved only 1 analog input. This analog signal comes from the DAC-8 (count starting from 0), which is located at "CH 1 OUT" SMA output in the "MONITORS" bin on the racks (attachment 2).
I have added another output channel to the DAF model both in software and in hardware. The DAF now also uses DAC-9 analog output which goes to the second analog input of the FiBox. The DAC-9 output is located at "CH 2 OUT" SMA output in the "MONITORS" bin on the racks (attachment 4).
After making the changes, the Fibox is shown in attacment 3.
Testing: The LSC input on passing through the DAF block is given through two different DAC outputs, to the same Fibox channel (one after the other), and the output is heard. More concrete testing will be done tomorrow. It will be as follows:
1) Currently, I need to search for a suitable cable that would connect the second channel of the output fibox to the audio mixer. After doing this, end to end testing of both channels will be done.
2) I could not access the AWG, probably because the DAQ was offline today afternoon. Using a signal from the AWG will give a more concrete testing of the stereo output.
3) After this, I will separate the two channels of the stereo completely (currectly they are seperated only at the DAF output stage)
4) I also will edit the medm gui appropriately.
I have added Enable buttons for each of the DSP blocks, and labels for the matrix elements. The input matrix takes inputs from each of the 4 channels: ADC1, ADC2, LSC and EXC, and routes them to the audio processing blocks (attachment 2). The output matrix (attachment 3) takes the outputs of the various DSP blocks and routes them to the output and then to the speakers.
Found PMC unlocked for many hours so I relocked it. IMC relocked by itself, but the input switch seems to be flickering to fast. Also the Keep Alive bit is not flashing.
I spent some time this afternoon reviving some of my CESAR/ESCOBAR shenanigans on the Y arm. I found it neccesary to adjust a few things.
Afterwards, ALSY noise levels were good.
I found an anti-aliasing circuit on the 40m wiki. It consists of A differential LPF made using THS4131 low noise differential op-amp (one of the main applications of which is preprocessing before the ADC), and a notch. I modified it to arrange for the desired bandwidth (about 8 kHz) and notch after the Nyquist frequency at 36 kHz. I simulated it to get the attached results:
Attachment 1: It shows the input PSD (same as the one posted in the previous elog), the filter transfer function, and The resulting output.
Attachment 2: The circuit schematic. The initial part using THS4131 is a differential LPF and the subsequent RC network is the notch.
Attachment 3: This shows the ratio of the aliased downconverted signal to the the in-band signal, representative of the contamination in each bin. Here too, the aliased signals are negligible as compared to the low frequencies but they are not negligible as compared to the higher frequencies (above 10 kHz) into which they would get downconverted due to sampling. However, here, the attenuation at 8kHz is less than 6 dB while in the previous circuit, it was about 12 dB. One problem with this circuit is at about 6kHz, there is aliased signal from the 65k to 98kHz band, but this can be taken care of by adding an LPF later.
Summary: The aim is to design an analog anti-aliasing (AA) filter placed before the ADC, whose function is to filter out components of the input spectrum that have frequencies higher than the Nyquist frequency. This needs to be done so that there is no contamination of aliased downconverted high-frequency signals into the ADC output. I have put down and simulated a circuit to do this, based on the spectra of a few interferometer signals that eric Provided. Attachment 1 shows such an input PSD, treated with whitening filter, before the AA. The sampling rate is 65536 Hz and hence the Nyquist freq. is 32768 Hz.
Motivation: Attachments 2 and 3 show the plot of required attenuation for various frequencies above the Nyquist. We can see a peak at 36 kHz, which will alias to about 29kHz. It will require about 70 dB attenuation here. This indicates that use of a notch filter combined with a low pass filter can be used.
Details of Schematic: Attachment 4 shows the schematic of a Boctor low pass notch filter, cascaded by a 2nd order LPF. The stopband frequency of the boctor filter can be tuned to around 36 kHz. Its main advantage for the boctor is better insensitivity to component value tolerances, use of a single op amp, and relatively independent tuning of parameters. The various component values are calculated from here. The transfer functions for the circuit shown in attachment 4 were simulated using TINA - a spice based simulation software. The transfer function is shown in attachment 5.
A few more calculations: Attachment 6 shows the output psd after the signal has been treated with AA. Attachments 7 and 8 show the ratio of aliased downconverted signal and the unaliased signal of the output. Here, we can see that above about 13 kHz, the ratios go above -40dB, which is apparently undesirable. However, we also see from the transfer function of the filter that the gain falls to less than -20dB after about this frequency, and the aliased signals are atleast 20 dB lower than this, atleast upto about 29 kHz in attachment 7 and about 25 kHz in attachment 8. This means that the aliased signals are negligible as compared to the low frequencies even if they are not negligible as compared to the higher frequencies (above 13 kHz) into which they would get downconverted due to sampling. But these higher frequencies (above 13 kHz) themselves are small.
The filter overall, is 4th order. Considering this and the above discussion, I need to decide what changes to make in the existing schematic. For now, I could discuss with eric to finalize the opamp and start building the pcb board design.
Below 100 Hz, I suppose this means that the X arm is now limited by the quadrature sum of the X and Y arm seismic noise.
I wiped down the cranes with wet towel and Mario our janitor did the chamber tops with the tubes.
The optical tabels were not touched.
The outside temp peaked at 44 C yesterday